OpenStudy (anonymous):

I don't understand the equation for a circle. (x - h)2 + (y - k)2 = r2 Can someone please explain it to me?

OpenStudy (anonymous):

(h,k) represents the center, and r represents the radius. To help your understanding, try coming up with a few simple circles and graphing them.

OpenStudy (anonymous):

I don't understand how to solve the equation when the numbers are put in

OpenStudy (anonymous):

For example, something like this: (x - 3)2 + (y - 7)2 = 25

OpenStudy (anonymous):

Basically, think of it like this. (h, k) is the center right? Since it says - h and - k in the circle equation, just bypass it and reverse the sign when putting it in or taking it out of the equation. Example: (x - 9)^2 + (y + 4)^2 = 9 The center is (9, -4) and the radius is 3. Just reverse the sign.

OpenStudy (anonymous):

This is the general equation for a circle, it isn't a specific circle. For a specific circle, the h, k, and r will all be defined. So you'll have, for example, \((x-1)^2+(y-2)^2=5^2.\) This would be a circle centered around (1,2) with radius 5. That just means that any x and y that make the equation true are on the circle.

OpenStudy (anonymous):

See how the signs are opposites?

OpenStudy (anonymous):

yeah that makes sense.